x^2d^2y/dx^2-xdy/dx+y=log x+sin(log x)+1/x


1
Expert's answer
2021-06-30T09:52:57-0400
"x^2y''-xy'+y=\ln x+\sin(\ln x)+\dfrac{1}{x}"

Let "x=e^t." Then


"y'=e^{-t}y'_t, y''=-e^{-2t}y'_t+e^{-2t}y''_{t^2}"

"\ln x=t"

"e^{2t}(-e^{-2t}y'_t+e^{-2t}y''_{t^2})-e^te^{-t}y'_t+y=t+\sin t+e^{-t}"

"y''_{t^2}-2y_t'+y=t+\sin t+e^{-t}"

Write the related homogeneous or complementary equation:


"y''_{t^2}-2y_t'+y=0"


The general solution of a nonhomogeneous equation is the sum of the general solution "y_h(x)" of the related homogeneous equation and a particular solution "y_p(x)" of the nonhomogeneous equation:


"y(x)=y_h(x)+y_p(x)"

Consider a homogeneous equation 


"y''_{t^2}-2y_t'+y=0"

Write the characteristic (auxiliary) equation:


"r^2-2r+1=0"

"(r-1)^2=0"

"r_1=1, r_2=1"

The general solution of the homogeneous equation is


"y_h(t)=C_1e^t+C_2te^t"


Let

"y_p=At+B+C\sin t+D\cos t+Ee^{-t}"

Then


"y_p'=A+C\cos t-D\sin t-Ee^{-t}"

"y_p''=-C\sin t-D\cos t+Ee^{-t}"

Substitute


"-C\sin t-D\cos t+Ee^{-t}"

"-2A+2C\cos t+2D\sin t+2Ee^{-t}"

"+At+B+C\sin t+D\cos t+Ee^{-t}"


"=t+\sin t+e^{-t}"

"A=1, B=2"




"C=0, D=\dfrac{1}{2}, E=\dfrac{1}{4}"

The general solution of a second order homogeneous differential equation be


"y(t)=C_1e^t+C_2te^t+t+2+\dfrac{1}{2}\cos t+\dfrac{1}{4}e^{-t}"




"y(x)=C_1x+C_2x\ln x+\ln x+2+\dfrac{1}{2}\cos (\ln x)+\dfrac{1}{4x}"